# On Identification of Dynamic Treatment Regimes with Proxies of Hidden Confounders

### v1

@@ 3.2 Main Result
There exist functions $h_{22}\left(\overline{y}_{2},\overline{w}_{2},\overline{a}_{2}\right)$ and $h_{21}\left(\overline{y}_{1},w_{1},\overline{a}_{2}\right)$ that satisfy
$f\left(Y_{2}=y_{2}|\overline{y}_{1},\overline{u}_{1},\overline{a}_{2}\right)=\sum_{\overline{w}_{2}}h_{22}\left(\overline{y}_{2},\overline{w}_{2},\overline{a}_{2}\right)f\left(\overline{w}_{2}|\overline{y}_{1},\overline{a}_{2},\overline{u}_{1}\right),$
and
$\sum_{\overline{w}_{2}}h_{22}\left(\overline{y}_{2},\overline{w}_{2},\overline{a}_{2}\right)f(y_{1}|y_{0},u_{0},a_{1})f\left(\overline{w}_{2}|\overline{y}_{1},a_{1},u_{0}\right)=\sum_{w_{1}}h_{21}\left(\overline{y}_{2},w_{1},\overline{a}_{2}\right)f\left(w_{1}|y_{0},a_{1},u_{0}\right),$
@@
The integral equation involving the $h_{22}$ function is similar to previous integral equations in the proxy literature (Miao et al., 2018; Ying et al., 2023; Tchetgen Tchetgen et al., 2024), with the left hand side being a conditional density given unmeasured confounders while the right hand side is a conditional mean function. Meanwhile, the integral equation involving $h_{21}$, to the best of our knowledge, has no precedent in the current literature. Rearranging the integral equation, we obtain
@@
Under Assumptions 1 and 2
$f\left(Y_{2}\left(a_{1},a_{2}\right)=y_{2},Y_{1}\left(a_{1}\right)=y_{1}|y_{0}\right)=\sum_{w_{1}}h_{21}\left(\overline{y}_{2},w_{1},\overline{a}_{2}\right)f\left(w_{1}|y_{0}\right)\text{ and }$ (2)
$f\left(Y_{1}\left(a_{1}\right)=y_{1}|y_{0}\right)=\sum_{y_{2}}\sum_{w_{1}}h_{21}\left(\overline{y}_{2},w_{1},\overline{a}_{2}\right)f\left(w_{1}|y_{0}\right).$ (3)
@@
For any square integrable $g_{1}$, we have that
${E}[g_{1}(\overline{U}_{1})\mid\overline{y}_{1},\overline{z}_{2},\overline{a}_{2}]=0\text{ implies }g_{1}(\overline{U}_{1})=0.$
In addition, for any square integrable $g_{0}$, we have that
${E}[g_{0}({U}_{0})\mid{y}_{0},{z}_{1},{a}_{1}]=0\text{ implies }g_{0}({U}_{0})=0.$
@@
Suppose there exist $h_{22}\left(\overline{y}_{2},\overline{w}_{2},\overline{a}_{2}\right)$ and $h_{21}\left(\overline{y}_{1},w_{1},\overline{a}_{2}\right)$ that satisfy
$f\left(Y_{2}=y_{2}|\overline{y}_{1},\overline{z}_{2},\overline{a}_{2}\right)$ $=\sum_{\overline{w}_{2}}h_{22}\left(\overline{y}_{2},\overline{w}_{2},\overline{a}_{2}\right)f\left(\overline{w}_{2}|\overline{y}_{1},\overline{a}_{2},\overline{z}_{2}\right)$ (4)
@@
respectively. If Assumption 3 also holds, then $h_{22}$ and $h_{21}$ also solve the latent integral equations from Assumption 2.
Lemma 1 gives a characterization of the bridge functions in terms of observable variables rather than the latent variables as in Assumption 2. Equipped with this result, identification of the optimal regime is possible if one can find $h_{22}\left(\overline{y}_{2},\overline{w}_{2},\overline{a}_{2}\right)$, $h_{21}\left(\overline{y}_{1},w_{1},\overline{a}_{2}\right)$, that satisfy
$f\left(Y_{2}=y_{2}|\overline{y}_{1},\overline{a}_{2},\overline{z}_{2}\right)$ $=\sum_{\overline{w}_{2}}h_{22}\left(\overline{y}_{2},\overline{w}_{2},\overline{a}_{2}\right)f\left(\overline{w}_{2}|\overline{y}_{1},\overline{a}_{2},\overline{z}_{2}\right),$
$\sum_{\overline{w}_{2}}h_{22}\left(\overline{y}_{2},\overline{w}_{2},\overline{a}_{2}\right)f\left(\overline{w}_{2}|\overline{y}_{1},a_{1},z_{1}\right)f(y_{1}|y_{0},\ a_{1},z_{1})$ $=\sum_{w_{1}}h_{21}\left(\overline{y}_{2},w_{1},\overline{a}_{2}\right)f\left(w_{1}|y_{0},a_{1},z_{1}\right).$
We note that uniqueness of solutions to the above equations is not necessary for identification of the optimal regime; any solutions suffice. We outline primitive conditions for the existence (and uniqueness) of solutions to integral equations in Section B of the supplementary material. Theorem 1 and Lemma 1 establish that the optimal rules/values at time points 2 and 1 are identified using only observed data as described in the following corollary.
@@
$d_{2}^{opt}\left(\overline{y}_{1},a_{1}\right)$ $=\underset{a_{2}}{\arg\max\ }{E}\left\{Y_{2}\left(a_{1},a_{2}\right)|\overline{Y}_{1}\left(a_{1}\right)=\overline{y}_{1}\right\}$
$=\underset{a_{2}}{\arg\max}\sum_{y_{2}}\frac{\sum_{w_{1}}h_{21}\left(\overline{y}_{2},w_{1},\overline{a}_{2}\right)f\left(w_{1}|y_{0}\right)}{\sum_{y_{2}}\sum_{w_{1}}h_{21}\left(\overline{y}_{2},w_{1},\overline{a}_{2}\right)f\left(w_{1}|y_{0}\right)}y_{2},$
$V_{2}^{opt}\left(a_{1},\overline{y}_{1}\right)$ $=\underset{a_{2}}{\max\ }{E}\left\{Y\left(a_{1},a_{2}\right)|\overline{Y}_{1}\left(a_{1}\right)=\overline{y}_{1}\right\}$
$=\underset{a_{2}}{\max}\sum_{y_{2}}\frac{\sum_{w_{1}}h_{21}\left(\overline{y}_{2},w_{1},\overline{a}_{2}\right)f\left(w_{1}|y_{0}\right)}{\sum_{y_{2}}\sum_{w_{1}}h_{21}\left(\overline{y}_{2},w_{1},\overline{a}_{2}\right)f\left(w_{1}|y_{0}\right)}y_{2},$
$d_{1}^{opt}\left(y_{0}\right)$ $=\underset{a_{1}}{\arg\max\ }{E}\left\{V_{2}^{opt}\left(a_{1},Y_{1}\left(a_{1}\right),Y_{0}\right)|Y_{0}=y_{0}\right\}$
$=\underset{a_{1}}{\arg\max}\sum_{y_{1}}\sum_{y_{2}}\sum_{w_{1}^{*}}h_{21}\left(\overline{y}_{2},w_{1}^{*},\overline{a}_{2}\right)f\left(w_{1}^{*}|y_{0}\right)V_{2}^{opt}\left(a_{1},\overline{y}_{1}\right).$
In principle, Corollary 1 can be used to estimate the optimal dynamic treatment regime. However, in the general case, the bridge functions can be challenging to estimate and compute nonparametrically as they may involve (conditional) densities. In particular, $h_{21}$’s nontrivial dependence on $h_{22}$ is distinct from other bridge functions in the literature (Ying et al., 2023; Tchetgen Tchetgen et al., 2024). In Section 4 below, we consider a setting only involving categorical variables as an initial crucial step towards inference in this challenging setting. Inference in the continuous case will be considered elsewhere given the significant challenges it presents.
@@ 3.3 A connection to the single time-point setting
In the single time-point setting, a single integral equation needs to be solved Miao et al. (2018), rather than nested ones. Specifically, Miao et al. (2018) require the existence of a function $h_{11}\left(w_{1},\overline{y}_{1},a_{1}\right)$ that solves the following so-called Fredholm integral equation of the first kind:
@@
4 Categorical Variables
4.1 Closed form expressions and estimation
@@
$=P(y_{2}\mid\overline{Z}_{2},\overline{y}_{1},\overline{a}_{2})P(\overline{W}_{2}\mid\overline{Z}_{2},\overline{y}_{1},\overline{a}_{2})^{-1}P(\overline{W}_{2},y_{1}\mid Z_{1},y_{0},a_{1})P(W_{1}\mid Z_{1},y_{0},a_{1})^{-1}P(W_{1}\mid y_{0}),\text{ and}$
$f(Y_{1}(a_{1})=y_{1}\mid y_{0})$
@@
Given the proceeding developments, simple substitution estimators are immediate. In fact, one may simply replace unknown population probabilities with empirical estimates. Then following the above, one obtains plug-in estimators $\widehat{f}(Y_{2}(a_{1},a_{2})=y_{2},Y_{1}(a_{1})=y_{1}\mid y_{0})$ and $\widehat{f}(Y_{1}(a_{1})=y_{1}\mid y_{0})$. Using these plug-in estimators, one then can find the treatment values maximizing corresponding value function (or minimizing regret) by appealing to Equation 1. In turn, one obtains estimates of the optimal treatment strategy at both time steps for any combination of histories of outcome and treatment variable. Stated explicitly:
$\widehat{d}_{2}^{proxy}\left(\overline{y}_{1},a_{1}\right)$
$\widehat{V}_{2}^{proxy}\left(a_{1},\overline{y}_{1}\right)$
With categorical variables, all (conditional) probabilities can be estimated from the observed cell counts.
@@ 5 Simulation results
5.1 Data generation and estimators
In this section, we investigate the finite sample performance of the estimator given in the previous section in the categorical confounding scenario. In a simulation study, we compare the plug-in proxy method to a naive approach that relies on the no unmeasured confounding assumption (NUCA), and an oracle procedure that has access to the true unmeasured confounder. More details about these methods can be found in Section E of the supplementary material. We simulate 11 binary variables according to a DAG that satisfies the conditional independence assumptions laid out in Section 1 from a saturated model. A DAG that encodes the temporal ordering and causal structure of variables is depicted in Figure 1. Exact details about the data generating mechanism are available in Section E of the supplementary material. The total number of iterations was 1000. We used saturated models and non-saturated log-linear models to fit the joint distribution of the observed data, from which all of the plug-in estimators can be derived respectively. For the log-linear models, we simply fit with higher order interactions up to some constant $K$, to deal with the possibility of sparse or empty cells. The observed data consists of 9 binary variables, so the fully saturated model corresponds to fitting a model with up to order 9 interactions. We compare NUCA with oracle, fully saturated proxy, and proxy with log-linear fit with 4,5,6,7,8-order interactions.
5.2 Results
First, in Table 1, we display, for each of the 4 sample sizes $N\in\{25000,50000,100000,250000\}$ and the methods oracle, NUCA, proxy, and proxy-LL-6, percentiles $q$ of the regrets, where regret is defined as the value of the optimal regime minus the value of the estimated optimal regime. The “proxy” method uses the fully saturated model and the “proxy-LL-6” method uses the log-linear model with order 6 interactions to fit the joint distribution of the observed data. Here, $q\in\{10,25,50,75,90\}$. Next, for the methods oracle, NUCA, and proxy, we plot the histograms of regrets for each sample size in Figure 2. Finally, we plot the difference in regret between NUCA and proxy in Figure 3. The results for the 4,5,7, and 8 way interactions are in the supplementary material. Since all data are binary, and the treatment decision at the first timepoint is based on one variable and the second based on three, there are a finite number $(1024=2^{2}\times 2^{8})$ of unique decision rules. Moreover, of these decision rules, there exist equivalence classes of regimes that induce the same exact value. This is because rules that are identical at the first time step will lead to only a subset of histories $(Y_{0},A_{1},Y_{1})$, and so rules identical at the first time step need only give the same decision on a subset of histories at time step 2 to yield the same value. The zeros in Table 1 indeed correspond to zero regret up to the machine precision in R, roughly 2.220446e-16.
@@ 5.2 Results
[…]
@@
[…]
Table 1: A summary of regrets for each method at varying sample sizes over 1000 iterations. The $q$ values denote the $q$th percentile of regret for the corresponding method.

### v2

@@ 3.2 Main Result
There exist functions $h_{22}(\overline{y}_{2},\overline{w}_{2},\overline{a}_{2})$ and $h_{21}(\overline{y}_{1},w_{1},\overline{a}_{2})$ that satisfy
$f(Y_{2}=y_{2}|\overline{y}_{1},\overline{u}_{1},\overline{a}_{2})=\sum_{\overline{w}_{2}}h_{22}(\overline{y}_{2},\overline{w}_{2},\overline{a}_{2})f(\overline{w}_{2}|\overline{y}_{1},\overline{a}_{2},\overline{u}_{1}),\text{ and }$
$\sum_{\overline{w}_{2}}h_{22}(\overline{y}_{2},\overline{w}_{2},\overline{a}_{2})f(y_{1}|y_{0},u_{0},a_{1})f(\overline{w}_{2}|\overline{y}_{1},a_{1},u_{0})=\sum_{w_{1}}h_{21}(\overline{y}_{2},w_{1},\overline{a}_{2})f(w_{1}\mid y_{0},a_{1},u_{0}),$
@@
The integral equation involving the $h_{22}$ function is similar to previous integral equations in the proxy literature (Miao et al., 2018; Ying et al., 2023; Tchetgen Tchetgen et al., 2024), with the left hand side being a conditional density given unmeasured confounders while the right hand side is a conditional mean function. Meanwhile, the integral equation involving $h_{21}$, does not resemble the typical conditional moment integral equations in the proxy literature. Rearranging the integral equation, we get
@@
Under Assumptions 1 and 2,
$f(Y_{2}(a_{1},a_{2})=y_{2},Y_{1}(a_{1})=y_{1}|y_{0})=\sum_{w_{1}}h_{21}(\overline{y}_{2},w_{1},\overline{a}_{2})f(w_{1}|y_{0})\text{ and }$
$f(Y_{1}(a_{1})=y_{1}|y_{0})=\sum_{y_{2}}\sum_{w_{1}}h_{21}(\overline{y}_{2},w_{1},\overline{a}_{2})f(w_{1}|y_{0}).$
@@
For any square integrable $g_{1}$, we have that ${E}\{g_{1}(\overline{U}_{1})\mid\overline{y}_{1},\overline{z}_{2},\overline{a}_{2}\}=0\text{ implies }g_{1}(\overline{U}_{1})=0$. In addition, for any square integrable $g_{0}$, we have that ${E}\{g_{0}({U}_{0})\mid{y}_{0},{z}_{1},{a}_{1}\}=0\text{ implies }g_{0}({U}_{0})=0.$
@@
Suppose Assumption 3 holds, and $h_{22}\left(\overline{y}_{2},\overline{w}_{2},\overline{a}_{2}\right)$ and $h_{21}\left(\overline{y}_{1},w_{1},\overline{a}_{2}\right)$ solve
$f\left(Y_{2}=y_{2}|\overline{y}_{1},\overline{z}_{2},\overline{a}_{2}\right)=\sum_{\overline{w}_{2}}h_{22}\left(\overline{y}_{2},\overline{w}_{2},\overline{a}_{2}\right)f\left(\overline{w}_{2}|\overline{y}_{1},\overline{a}_{2},\overline{z}_{2}\right),$ (3)
@@
respectively. Then $h_{22}$ and $h_{21}$ also solve the latent integral equations from Assumption 2.
Lemma 1 gives a characterization of the bridge functions in terms of observable variables rather than the latent variables as in Assumption 2. Equipped with this result, identification of the optimal regime is possible if one can find $h_{22}\left(\overline{y}_{2},\overline{w}_{2},\overline{a}_{2}\right)$, $h_{21}\left(\overline{y}_{1},w_{1},\overline{a}_{2}\right)$, that respectively satisfy the equations from Equation 3. We note that uniqueness of solutions to the above equations is not necessary for identification of the optimal regime; any solutions suffice. We outline primitive conditions for the existence (and uniqueness) of solutions to integral equations in Section B of the supplementary material. Theorem 1 and Lemma 1 establish that the optimal rules/values at time points 2 and 1 are identified using only observed data as described in the following corollary.
@@
$d_{2}^{opt}\left(\overline{y}_{1},a_{1}\right)$ $=\underset{a_{2}}{\arg\max}\sum_{y_{2}}\frac{\sum_{w_{1}}h_{21}\left(\overline{y}_{2},w_{1},\overline{a}_{2}\right)f\left(w_{1}|y_{0}\right)}{\sum_{y_{2}}\sum_{w_{1}}h_{21}\left(\overline{y}_{2},w_{1},\overline{a}_{2}\right)f\left(w_{1}|y_{0}\right)}y_{2},$
$V_{2}^{opt}\left(a_{1},\overline{y}_{1}\right)$ $=\underset{a_{2}}{\max}\sum_{y_{2}}\frac{\sum_{w_{1}}h_{21}\left(\overline{y}_{2},w_{1},\overline{a}_{2}\right)f\left(w_{1}|y_{0}\right)}{\sum_{y_{2}}\sum_{w_{1}}h_{21}\left(\overline{y}_{2},w_{1},\overline{a}_{2}\right)f\left(w_{1}|y_{0}\right)}y_{2},$
$d_{1}^{opt}\left(y_{0}\right)$ $=\underset{a_{1}}{\arg\max}\sum_{y_{1}}\sum_{y_{2}}\sum_{w_{1}^{*}}h_{21}\left(\overline{y}_{2},w_{1}^{*},\overline{a}_{2}\right)f\left(w_{1}^{*}|y_{0}\right)V_{2}^{opt}\left(a_{1},\overline{y}_{1}\right).$
In principle, Corollary 1 can be used to estimate the optimal dynamic treatment regime. However, in the general case, the bridge functions can be challenging to estimate and compute nonparametrically as they may involve (conditional) densities. In particular, $h_{21}$’s nontrivial dependence on $h_{22}$ is distinct from other bridge functions in the literature (Ying et al., 2023; Tchetgen Tchetgen et al., 2024). In Section 4.1 below, we consider a setting only involving categorical variables as an initial crucial step towards inference in this challenging setting. Inference in the continuous case will be considered elsewhere given the significant challenges it presents.
@@ 3.3 A connection to the single time-point setting
In the single time-point setting, a single integral equation needs to be solved, rather than nested ones. Specifically, Miao et al. (2018) require the existence of a function $h_{11}\left(w_{1},\overline{y}_{1},a_{1}\right)$ that solves the following so-called Fredholm integral equation of the first kind:
@@
4 Estimation of Bridge Functions
4.1 Categorical Variables
@@
$f(Y_{1}(a_{1})=y_{1}\mid y_{0})=P(y_{1}\mid\overline{Z}_{1},\overline{y}_{0},\overline{a}_{1})P(W_{1}\mid Z_{1},y_{0},a_{1})^{-1}P(W_{1}\mid y_{0}).$
@@
Given the proceeding developments, simple substitution estimators are immediate. In fact, one may simply replace unknown population probabilities with empirical estimates. Then following the above, one obtains plug-in estimators $\widehat{f}(Y_{2}(a_{1},a_{2})=y_{2},Y_{1}(a_{1})=y_{1}\mid y_{0})$ and $\widehat{f}(Y_{1}(a_{1})=y_{1}\mid y_{0})$. Using these plug-in estimators, one then can find the treatment values maximizing corresponding value function (or minimizing regret) by appealing to Equation 2. In turn, one obtains estimates of the optimal treatment strategy at both time steps for any combination of histories of outcome and treatment variable. Explicit expressions are contained in Section E of the supplementary material. With categorical variables, all (conditional) probabilities can be estimated from the observed cell counts.
4.2 Categorical Outcomes
In this subsection, we outline another situation in which it is feasible to estimate solutions to the integral equations. As noted previously, the integral equations described above are extremely difficult to estimate in the general case, since they involve conditional densities. Therefore, we consider a simpler case where $Y_{1}$ and $Y_{2}$ are assumed to be discrete, but all other variables are unrestricted. In such a case, the first observed bridge function from Equation 3 can be characterized as follows:
$E[\mathbf{1}(Y_{2}=y_{2})\mid\overline{y}_{1},\overline{a}_{2},\overline{z}_{2}]$ $=\sum_{\overline{w}_{2}}h_{22}\left(\overline{y}_{2},\overline{w}_{2},\overline{a}_{2}\right)f\left(\overline{w}_{2}|\overline{y}_{1},\overline{a}_{2},\overline{z}_{2}\right)$ (5)
The second bridge function can be rewritten as follows:
$\sum_{\overline{w}_{2}}h_{22}\left(\overline{y}_{2},\overline{w}_{2},\overline{a}_{2}\right)f\left(\overline{w}_{2}|\overline{y}_{1},a_{1},z_{1}\right)f(y_{1}|y_{0},\ a_{1},z_{1})$ $=\sum_{w_{1}}h_{21}\left(\overline{y}_{2},w_{1},\overline{a}_{2}\right)f\left(w_{1}|y_{0},a_{1},z_{1}\right)\iff$ (6)
$\sum_{\overline{w}_{2}}h_{22}\left(\overline{y}_{2},\overline{w}_{2},\overline{a}_{2}\right)f\left(\overline{w}_{2},y_{1}|{y}_{0},a_{1},z_{1}\right)$ $=\sum_{w_{1}}h_{21}\left(\overline{y}_{2},w_{1},\overline{a}_{2}\right)f\left(w_{1}|y_{0},a_{1},z_{1}\right)\iff$
$\sum_{\overline{w}_{2},\tilde{y}_{1}}h_{22}\left(\overline{y}_{2},\overline{w}_{2},\overline{a}_{2}\right)\mathbf{1}(\tilde{y_{1}}=y_{1})f\left(\overline{w}_{2},\tilde{y}_{1}|{y}_{0},a_{1},z_{1}\right)$ $=\sum_{w_{1}}h_{21}\left(\overline{y}_{2},w_{1},\overline{a}_{2}\right)f\left(w_{1}|y_{0},a_{1},z_{1}\right)\iff$
$E_{\overline{W}_{2},{Y}_{1}}[h_{22}\left(\overline{y}_{2},\overline{W}_{2},\overline{a}_{2}\right)\mathbf{1}({Y_{1}}=y_{1})\mid{y}_{0},a_{1},z_{1}]$ $=E_{W_{1}}[h_{21}\left(\overline{y}_{2},W_{1},\overline{a}_{2}\right)\mid y_{0},a_{1},z_{1}].$
These characterizations imply that for fixed $\overline{a}_{2},y_{1},y_{2}$, the corresponding $h_{22}$ and $h_{21}$ bridge function characterizations are conditional moment equations. Min-max estimators for such bridge functions have been proposed and applied in earlier work (Dikkala et al., 2020; Ghassami et al., 2022).
4.3 Categorical Outcomes - Treatment Bridge Function
When the outcomes $Y_{1}$ and $Y_{2}$ are discrete, we can also identify the identify the joint distribution of the discrete potential outcomes using different bridge functions, which we will refer to as treatment bridge functions. Explicitly, we assume the following:
Assumption 5 (Time-varying Treatment Bridge Functions).
There exist functions $q_{11}\left({y}_{0},{z}_{1},{a}_{1}\right)$ and $q_{12}\left(\overline{y}_{1},\overline{z}_{2},\overline{a}_{2}\right)$ that satisfy
$\frac{1}{P(A_{1}=a_{1}\mid u_{0},y_{0})}=\sum_{z_{1}}q_{11}\left({y}_{0},{z}_{1},{a}_{1}\right)f(z_{1}\mid a_{1},u_{0},y_{0}),$
and
$\frac{E[q_{11}\left({y}_{0},{Z}_{1},{a}_{1}\right)\mid a_{1},u_{0},y_{0}]}{P(A_{2}=a_{2}\mid a_{1},\overline{u}_{1},\overline{y}_{1})}=\sum_{\overline{z}_{2}}q_{12}\left(\overline{y}_{1},\overline{z}_{2},\overline{a}_{2}\right)f(\overline{z}_{2}\mid\overline{y}_{1},\overline{u}_{1},\overline{a}_{2}),$
respectively.
The treatment bridge functions can be used to identify the joint distribution of the potential outcomes, as described in the following theorem:
Theorem 2.
Suppose outcomes $Y_{2}$ and $Y_{1}$ are discrete. Under Assumptions 1 and 5,
$f\left(Y_{2}\left(a_{1},a_{2}\right)=y_{2},Y_{1}\left(a_{1}\right)=y_{1}|y_{0}\right)=E[q_{12}\left(\overline{y}_{1},\overline{Z}_{2},\overline{a}_{2}\right)\mathbf{1}\{Y_{2}=y_{2},Y_{1}=y_{1}\}\mathbf{1}\{\overline{A}_{2}=\overline{a}_{2}\}\mid y_{0}]\text{ and }$
$f\left(Y_{1}\left(a_{1}\right)=y_{1}|y_{0}\right)=\sum_{y_{2}}E[q_{12}\left(\overline{y}_{1},\overline{Z}_{2},\overline{a}_{2}\right)\mathbf{1}\{Y_{2}=y_{2},Y_{1}=y_{1}\}\mathbf{1}\{\overline{A}_{2}=\overline{a}_{2}\}\mid y_{0}],$
where $\mathbf{1}$ denotes the indicator function.
We now introduce an analogous condition to Assumption 3 that will characterize the treatment bridge functions in terms of the observed data.
Assumption 6 (W-Completeness/Relevance).
For any square integrable $g_{1}$, we have that
${E}[g_{1}(\overline{U}_{1})\mid\overline{y}_{1},\overline{w}_{2},\overline{a}_{2}]=0\text{ implies }g_{1}(\overline{U}_{1})=0.$
In addition, for any square integrable $g_{0}$, we have that
${E}[g_{0}({U}_{0})\mid{y}_{0},{w}_{1},{a}_{1}]=0\text{ implies }g_{0}({U}_{0})=0.$
Lemma 2.
Suppose there exist $q_{12}\left(\overline{y}_{2},\overline{z}_{2},\overline{a}_{2}\right)$ and $q_{11}\left(\overline{y}_{1},z_{1},\overline{a}_{2}\right)$ that satisfy
$\frac{1}{P(A_{1}=a_{1}\mid w_{1},y_{0})}=\sum_{z_{1}}q_{11}\left({y}_{0},{z}_{1},{a}_{1}\right)f(z_{1}\mid a_{1},w_{1},y_{0},)$ (7)
$\frac{E[q_{11}\left({y}_{0},{Z}_{1},{a}_{1}\right)\mid a_{1},w_{1},y_{0}]}{P(A_{2}=a_{2}\mid a_{1},\overline{w}_{2},\overline{y}_{1})}=\sum_{\overline{z}_{2}}q_{12}\left(\overline{y}_{1},\overline{z}_{2},\overline{a}_{2}\right)f(\overline{z}_{2}\mid\overline{y}_{1},\overline{w}_{2},\overline{a}_{2}),$
respectively. If Assumption 6 also holds, then $q_{11}$ and $q_{12}$ also solve the latent integral equations from Assumption 5.
Similar to Lemma 1, Lemma 2 gives a characterization of the treatment bridge functions in terms of observable variables rather than the latent variables. The equations in (7) are conditional moment equations, and so the min-max estimators from Dikkala et al. (2020) can be readily used to estimate the treatment bridge functions.
@@ 5 Simulation results
In this section, we investigate the finite sample performance of the closed-form estimator given in the previous section in the setting where all variables are categorical . In a simulation study, we compare the plug-in proxy method to a naive approach that relies on the no unmeasured confounding assumption (NUCA), and an oracle procedure that has access to the true unmeasured confounder. More details about these methods can be found in Section E of the supplementary material. We simulate 11 binary variables according to the DAG in Figure 1 that satisfies the conditional independence assumptions laid out in Section 1 from a saturated model. Exact details about the data generating mechanism are available in Section E of the supplementary material. The total number of iterations was 1000. We used saturated models and non-saturated log-linear models to fit the joint distribution of the observed data, from which all of the plug-in estimators can be derived respectively. For the log-linear models, we simply fit with higher order interactions up to some constant $K$, to deal with the possibility of sparse or empty cells. The observed data consists of 9 binary variables, so the fully saturated model corresponds to fitting a model with up to order 9 interactions. We compare NUCA with oracle, fully saturated proxy, and proxy with log-linear fit with 4,5,6,7,8-order interactions. First, in Table 1, we display, for each of the 4 sample sizes $N\in\{25000,50000,100000,250000\}$ and the methods oracle, NUCA, proxy, and proxy-LL-6, percentiles $q$ of the regrets (the optimal regime value minus the estimated regime value). The “proxy” method uses the fully saturated model and the “proxy-LL-6” method uses the log-linear model with order 6 interactions to fit the joint distribution of the observed data. Here, $q\in\{10,25,50,75,90\}$. More detailed results and results for the 4,5,7, and 8 way interactions are in the supplementary material. Since all data are binary, and the treatment decision at the first timepoint is based on one variable and the second based on three, there are a finite number of unique decision rules. Moreover, there exist equivalence classes of regimes that induce the same exact value. This is because rules that are identical at the first time step will lead to only a subset of histories $(Y_{0},A_{1},Y_{1})$, and so rules identical at the first time step need only give the same decision on a subset of histories at time step 2 to yield the same value. The $\epsilon$ in Table 1 correspond to zero regret up to the machine precision in R, roughly ϵ=italic-$\epsilon=$ 2.220446e-16.
@@
25,000 proxy $<\epsilon$ $<\epsilon$ 0.01503 0.05214 0.09824 0.03705
NUCA $<\epsilon$ $<\epsilon$ 0.09159 0.09159 0.09159 0.06420
oracle $<\epsilon$ $<\epsilon$ $<\epsilon$ $<\epsilon$ $<\epsilon$ 0.00001
proxy-LL-6 $<\epsilon$ $<\epsilon$ 0.01503 0.04746 0.09824 0.03574
50,000 proxy $<\epsilon$ $<\epsilon$ 0.00932 0.03711 0.09159 0.02588
NUCA $<\epsilon$ 0.09159 0.09159 0.09159 0.09159 0.07052
oracle $<\epsilon$ $<\epsilon$ $<\epsilon$ $<\epsilon$ $<\epsilon$ $<\epsilon$
proxy-LL-6 $<\epsilon$ $<\epsilon$ $<\epsilon$ 0.03711 0.09159 0.02626
100,000 proxy $<\epsilon$ $<\epsilon$ $<\epsilon$ 0.01503 0.09159 0.01894
NUCA $<\epsilon$ 0.09159 0.09159 0.09159 0.09159 0.07922
oracle $<\epsilon$ $<\epsilon$ $<\epsilon$ $<\epsilon$ $<\epsilon$ $<\epsilon$
proxy-LL-6 $<\epsilon$ $<\epsilon$ $<\epsilon$ 0.01503 0.09159 0.01927
250,000 proxy $<\epsilon$ $<\epsilon$ $<\epsilon$ 0.01503 0.06117 0.01521
@@
oracle $<\epsilon$ $<\epsilon$ $<\epsilon$ $<\epsilon$ $<\epsilon$ $<\epsilon$
proxy-LL-6 $<\epsilon$ $<\epsilon$ $<\epsilon$ 0.01503 0.04613 0.01364
Table 1: A summary of regret for each method at varying sample sizes over 1000 iterations. $q$ denotes the $q$th percentile of regret.